OpenStudy (anonymous):

Find the slope of the line that passes through the points (5, 8) and (9, 4). A. 3 B. 1 C. -1 D. 4/3

jimthompson5910 (jim_thompson5910):

Use the slope formula m = (y2-y1)/(x2-x1) m = ( 4- 8)/(9-5) I'll let you finish

OpenStudy (anonymous):

-1x + 13 so -1